# Sequential model confidence sets

### v3

@@ 3.3 SMCSs for the weakly superior objects
Our SMCS construction for the weakly superior objects defined at (3) builds on time-uniform confidence regions. Proofs for this section are given in Supplement E.
@@
^$\widehat{\mathcal{M}}_{t}=\left\{i\in\mathcal{M}_{0}\mid C_{t}\cap\{\boldsymbol{X}\in\mathbb{R}_{0}^{m\times m}\mid x_{ij}\leq 0\}\not=\emptyset\text{ for all }j\neq i\right\},\quad t\in\mathbb{N}.$ (9)
@@
For any $\alpha\in(0,1)$, the sequence (^$(\widehat{\mathcal{M}}_{t})_{t\in\mathbb{N}}$ defined at (9) is an SMCS at level $\alpha$ for the weakly superior objects $(\mathcal{M}^{\textrm{w},\star}_{t})_{t\in\mathbb{N}}$. Its running intersection ~$\widetilde{\mathcal{M}}_{t}=\bigcap_{r\leq t}\widehat{\mathcal{M}}_{r},t\in\mathbb{N}$ is an SMCS at level $\alpha$ for the uniformly weakly superior objects $\mathcal{M}^{\textrm{uw},\star}$.
@@
Under the conditions of Proposition 3.5, the definition of ^$\widehat{\mathcal{M}}_{t}$ at (9) simplifies:
^$\widehat{\mathcal{M}}_{t}=\left\{i\in\mathcal{M}_{0}\mid C_{t}\cap\{\boldsymbol{X}\in\mathbb{R}_{0}^{m\times m}\mid x_{ij}=0\}\not=\emptyset\text{for all }j\neq i\right\},\quad t\in\mathbb{N}.$ (10)
@@ 3.4 Predictable bounds and betting schemes
Assume that $\lvert d_{ij,t}\rvert\leq c_{ij,t}/2$ for all $t\in\mathbb{N}$ for some predictable $(c_{ij,t})_{t\in\mathbb{N}}$. Then, the transformed loss differences $\tilde{d}_{ij,t}=d_{ij,t}/c_{ij,t},t\in\mathbb{N}$ are uniformly bounded with constant $c=1$ that is $\lvert\tilde{d}_{ij,t}\rvert\leq 1/2$ for all $t\in\mathbb{N}$. However, if we use the transformed loss differences, we target (possibly) different sequences of superior objects ~$\tilde{\mathcal{M}}^{\textrm{uw},\star}=\{i\in\mathcal{M}_{0}\mid\tilde{\Delta}_{ij,t}\leq 0,\text{for all}\;j\in\mathcal{M}_{0},t\in\mathbb{N}\}$ and ~$\tilde{\mathcal{M}}^{\textrm{w},\star}_{t}=\{i\in\mathcal{M}_{0}\mid\tilde{\Delta}_{ij,t}\leq 0,\text{for all}\;j\in\mathcal{M}_{0}\}$ for $\tilde{\Delta}_{ij,t}=(1/t)\sum_{r=1}^{t}\mathbb{E}(\tilde{d}_{ij,r}\mid\mathcal{F}_{r-1})=(1/t)\sum_{r=1}^{t}\mu_{ij,r}/c_{ij,r}.$
In Section 5, we use the given transformation to convert conditionally bounded CRPS differences of wind gust forecasts into uniformly bounded loss differences. Importantly, for any proper scoring rule (consistent scoring function) ${\rm S}$, the scaled function $\tilde{{\rm S}}={\rm S}/c$ is proper (consistent) as well for any $c>0$, an observation which justifies the transformation from a theoretical perspective. This property can be extended in the sequential setting, see Supplement A. With the given scaling, all observations have the same maximal impact on the ranking, an effect which seems appealing for many (but surely not for all) applications.
@@
For important choices of loss functions such as the quadratic or logarithmic score, the corresponding loss differences are not conditionally bounded. However, if the loss differences emerge from (or may be bounded in the tails by) a parametric family of distributions, there are other well-studied e-processes available; see, e.g., Howard et al. (2020). In Supplement F.1, we provide a simulation example of mean forecasts with Gaussian errors. The resulting score differences are unbounded but still sub-exponential, which allows to use the e-processes of Proposition 3.3.
@@ 4.1 Simulation 1
We ran Simulation 1 also for the original MCS by Hansen et al. (2011). Although its size is generally better than that of the SMCSs, the coverage rate drops below the nominal level $0.91-\alpha=$ 0.9 to about $0.8$. The sequential implementation of the MCS is documented in Supplement F.2, along with another simulation in a more challenging, non-Gaussian setting. There, the coverage rates of the MCS are heavily compromised, in contrast to those of the SMCS.
@@ 4.2 Simulation 2
It could be interesting to monitor the e-processes that we use to collect evidence on predictive performance over time. Due to the large number of models considered here, this is not so practical, but we illustrate this idea in Supplement F.1, where only nine models are considered in the experiment.
@@ 4.3 Simulation 3
We consider the simulation setting from Section 2.1. Let $i=1,2,3$ be forecasters issuing median predictions $m_{i,t}=Y_{t}+\varepsilon_{i,t}$, $t\in\mathbb{N}$, for i.i.d. standard normally distributed $(Y_{t})_{t\in\mathbb{N}}$ and $\varepsilon_{1,t}=\beta,\varepsilon_{2,t}=\gamma^{t},\varepsilon_{3,t}=\delta t,$ for $\beta,\delta>0$ and $10<\gamma<$ 1. For the loss $L(m,y)=0.5(\Phi(m)-\Phi(y))$, we get by independence that $2\mu_{ij,t}=\mathbb{E}(\Phi(m_{i,t}))-\mathbb{E}(\Phi(m_{j,t}))=\Phi\left(\varepsilon_{i,t}/\sqrt{2}\right)-\Phi\left(\varepsilon_{j,t}/\sqrt{2}\right)$ for $t\in\mathbb{N}$, $i,j=1,2,3$, thus
@@
We let $\beta=0.6,\gamma=0.998$ and $\delta=0.008$. Then, one can numerically show that $\mathcal{M}^{\textrm{w},\star}_{t}=\{3\}$ for $t\leq 153$, $\mathcal{M}^{\textrm{w},\star}_{t}=\{1\}$ for $550153<t<$ 550, and $\mathcal{M}^{\textrm{w},\star}_{t}=\{2\}$ for $t\geq 550$.
@@ 5.1.3 Results
Throughout the paper, we have used a significance level equal to $\alpha=0.1$, but clearly, there is freedom to do otherwise. In particular, we could hope to detect some defective models even earlier by increasing $\alpha$ if we are willing to increase the type-I error control level.
@@ 6 Discussion
We have provided coverage guarantees of all superior models with high probability uniformly over time. Other coverage guarantees could be of interest but require somewhat different approaches. The literature on inference on the argmin of a vector of means typically focuses on the guarantee that each superior model is covered with high probability. Future work could consider time-uniform extensions of such guarantees. Furthermore, one could aim for a time-uniform bound of the expected fraction of wrongly rejected models by the total number of rejected models. This can be achieved by time-uniform control of the false discovery rate and we give first results in this direction in Supplement D.
We have mainly focused on (conditionally) bounded score differences. However, important loss functions such as the squared error or the logarithmic score do not satisfy this requirement. Nevertheless, our general theory is not limited to bounded score differences. We have provided a simulation study for SMCSs with the squared error loss in Supplement F.1. Nonetheless, further research is warranted to extend the understanding and implementation of SMCSs to encompass more general loss functions. Finally, it would be interesting to examine how our results extend to important information criteria, and to apply sequential model confidence sets for sequential model selection.
@@ Acknowledgments
The comments of three anonymous referees significantly improved the quality of the paper. We thank Lutz Dümbgen for valuable discussions and inputs. Sebastian Arnold and Johanna Ziegel gratefully acknowledge financial support from the Swiss National Science Foundation. Computations for the simulations and parts of the wind gust case study have been performed on UBELIX (https://ubelix.unibe.ch/), the HPC cluster of the University of Bern. Benedikt Schulz gratefully acknowledges funding within the project C5 “Dynamical feature-based ensemble postprocessing of wind gusts within European winter storms” of the Transregional Collaborative Research Center SFB/TRR 165 “Waves to Weather” funded by the German Research Foundation (DFG). We thank Cristina Primo, Reinhold Hess and Sebastian Trepte for providing the forecast and observation data, and Robert Redl for assistance in data handling.
@@ Appendix B Conditionally bounded score differences
[…]
Proof.
We have $\delta(y)=\int F_{1}(x)^{2}-F_{2}(x)^{2}\,\mathrm{d}x+2\int_{y}^{\infty}F_{2}(x)-F_{1}(x)\,\mathrm{d}x,$ where the first summand is independent of $y\in\mathbb{R}$. The function $x\mapsto F_{2}(x)-F_{1}(x)$ is bounded in $[-1,1]$, with limit 00 as $x\to\pm\infty$ and changes sign at the crossing points of $F_{1}$ and $F_{2}$. ∎
Lemma B.2.
@@
In the spirit of Vovk and Wang (2021), we say that e-processes $(E_{1,t}^{\star})_{t\in\mathbb{N}},\dots,(E_{m,t}^{\star})_{t\in\mathbb{N}}$ for the hypotheses $\mathcal{H}_{1},\dots,\mathcal{H}_{m}$ are family-wise valid (FWV) if there exists a dominating family of e-processes, that is, if there exists a family $\{(E^{\mathbb{Q}}_{t})_{t\in\mathbb{N}}\mid\mathbb{Q}\in\mathfrak{B}(\Omega)\}$ such that $(E^{\mathbb{Q}}_{t})_{t\in\mathbb{N}}$ is an e-process for ${}\{\mathbb{Q}\}$ for all $\mathbb{Q}\in\mathfrak{B}(\Omega)$, and for any $i=1,\dots,m$ with $\mathbb{Q}\in\mathcal{H}_{i}$, it holds that $E_{t}^{\mathbb{Q}}\geq E_{i,t}^{\star}$.
@@
Suppose that we adjust the e-processes $(E_{1,t})_{t\in\mathbb{N}},\dots,(E_{m,t})_{t\in\mathbb{N}}$ by the closure principle with respect to the same e-merging function $f$ for all $t\in\mathbb{N}$. Then it follows that $(E^{\star}_{1,t})_{t\in\mathbb{N}},\dots,(E^{\star}_{m,t})_{t\in\mathbb{N}}$ are family-wise valid e-processes for the hypotheses $\mathcal{H}_{1},\dots,\mathcal{H}_{m}$, since $(E_{t}^{\mathbb{Q}})_{t\in\mathbb{N}}$ are e-processes. To see this, let $\tau$ be a stopping time. Then, $\mathbb{E}_{\mathbb{Q}}(E_{\tau}^{\mathbb{Q}})=\mathbb{E}_{\mathbb{Q}}(f(E_{i,\tau},i\in I_{\mathbb{Q}}))\leq 1$, since $E_{i,\tau}$, $i\in I_{\mathbb{Q}}$ are e-values under $\mathbb{Q}$.
@@
where $E_{t}^{\mathbb{Q}}=f(E_{i\cdot,t},i\in I_{\mathbb{Q}})$ with $I_{\mathbb{Q}}=\{i\mid\mathbb{Q}\in\mathcal{H}^{\bullet}_{i\cdot}\}$. It follows directly that the running intersection ~$\widetilde{\mathcal{M}}_{t}=\bigcap_{r\leq t}\widehat{\mathcal{M}}_{r},t\in\mathbb{N}$, is an SMCS as well.∎
@@
^$\widehat{\mathcal{M}}_{t}=\left\{i\in\mathcal{M}_{0}\mid E_{ij,t}^{\star}<1/\alpha,\text{ for all }j\in\mathcal{M}_{0}\right\},\quad t\in\mathbb{N}.$
@@
Appendix D SMCSs by FDR control

### v4

@@ 3.3 SMCSs for the weakly superior objects
Our SMCS construction for the weakly superior objects defined at (3) builds on time-uniform confidence regions. Proofs for this section are given in Supplement F.
@@
$\widehat{\mathcal{M}}_{t}=\left\{i\in\mathcal{M}_{0}\mid C_{t}\cap\{\boldsymbol{X}\in\mathbb{R}_{0}^{m\times m}\mid x_{ij}\leq 0\}\not=\emptyset\text{ for all }j\neq i\right\},\quad t\in\mathbb{N}.$ (9)
@@
For any $\alpha\in(0,1)$, the sequence $(\widehat{\mathcal{M}}_{t})_{t\in\mathbb{N}}$ defined at (9) is an SMCS at level $\alpha$ for the weakly superior objects $(\mathcal{M}^{\textrm{w},\star}_{t})_{t\in\mathbb{N}}$. Its running intersection $\widetilde{\mathcal{M}}_{t}=\bigcap_{r\leq t}\widehat{\mathcal{M}}_{r},t\in\mathbb{N}$ is an SMCS at level $\alpha$ for the uniformly weakly superior objects $\mathcal{M}^{\textrm{uw},\star}$.
Similarly to Remark 2, we may construct more powerful SMCSs if we target a suitable marginal coverage guarantee, see Supplement D for details.
@@
Under the conditions of Proposition 3.5, the definition of $\widehat{\mathcal{M}}_{t}$ at (9) simplifies:
$\widehat{\mathcal{M}}_{t}=\left\{i\in\mathcal{M}_{0}\mid C_{t}\cap\{\boldsymbol{X}\in\mathbb{R}_{0}^{m\times m}\mid x_{ij}=0\}\not=\emptyset\text{ for all }j\neq i\right\},\quad t\in\mathbb{N}.$ (10)
@@ 3.4 Predictable bounds and betting schemes
Assume that $\lvert d_{ij,t}\rvert\leq c_{ij,t}/2$ for all $t\in\mathbb{N}$ for some predictable $(c_{ij,t})_{t\in\mathbb{N}}$. Then, the transformed loss differences $\tilde{d}_{ij,t}=d_{ij,t}/c_{ij,t},t\in\mathbb{N}$ are uniformly bounded with constant $c=1$ that is $\lvert\tilde{d}_{ij,t}\rvert\leq 1/2$ for all $t\in\mathbb{N}$. However, if we use the transformed loss differences, we target (possibly) different sequences of superior objects $\tilde{\mathcal{M}}^{\textrm{uw},\star}=\{i\in\mathcal{M}_{0}\mid\tilde{\Delta}_{ij,t}\leq 0,\text{for all}\;j\in\mathcal{M}_{0},t\in\mathbb{N}\}$ and $\tilde{\mathcal{M}}^{\textrm{w},\star}_{t}=\{i\in\mathcal{M}_{0}\mid\tilde{\Delta}_{ij,t}\leq 0,\text{for all}\;j\in\mathcal{M}_{0}\}$ for $\tilde{\Delta}_{ij,t}=(1/t)\sum_{r=1}^{t}\mathbb{E}(\tilde{d}_{ij,r}\mid\mathcal{F}_{r-1})=(1/t)\sum_{r=1}^{t}\mu_{ij,r}/c_{ij,r}.$
In Section 5, we use the given transformation to convert conditionally bounded CRPS differences of wind gust forecasts into uniformly bounded loss differences. Computational feasibility of this approach is due to the forecasts being either parametric or with finite support, see Supplement B for details. Importantly, for any proper scoring rule (consistent scoring function) $\rm S$}, the scaled function $\tilde{{\rm S}}={\rm S}/c$ is proper (consistent) as well for any $c>0$, an observation which justifies the transformation from a theoretical perspective. This property can be extended in the sequential setting, see Supplement A. With the given scaling, all observations have the same maximal impact on the ranking, an effect which seems appealing for many (but surely not for all) applications.
@@
For important choices of loss functions such as the quadratic or logarithmic score, the corresponding loss differences are not conditionally bounded. However, if the loss differences emerge from (or may be bounded in the tails by) a parametric family of distributions, there are other well-studied e-processes available; see, e.g., Howard et al. (2020). In Supplement G.1, we provide a simulation example of mean forecasts with Gaussian errors. The resulting score differences are unbounded but still sub-exponential, which allows to use the e-processes of Proposition 3.3.
@@ 4.1 Simulation 1
We ran Simulation 1 also for the original MCS by Hansen et al. (2011). Although its size is generally better than that of the SMCSs, the coverage rate drops below the nominal level $1-\alpha=0.9$ to about $0.8$. The sequential implementation of the MCS is documented in Supplement G.2, along with another simulation in a more challenging, non-Gaussian setting. There, the coverage rates of the MCS are heavily compromised, in contrast to those of the SMCS.
@@ 4.2 Simulation 2
It could be interesting to monitor the e-processes that we use to collect evidence on predictive performance over time. Due to the large number of models considered here, this is not so practical, but we illustrate this idea in Supplement G.1, where only nine models are considered in the experiment.
In Supplement D, we illustrate the potential gain in power by replacing simultaneous coverage by marginal coverage (see Remark 2) by comparing the average size of the SMCS in Simulations 1 and 2 with and without multiple testing correction.
@@ 4.3 Simulation 3
We consider the simulation setting from Section 2.1. Let $i=1,2,3$ be forecasters issuing median predictions $m_{i,t}=Y_{t}+\varepsilon_{i,t}$, $t\in\mathbb{N}$, for i.i.d. standard normally distributed $(Y_{t})_{t\in\mathbb{N}}$ and $\varepsilon_{1,t}=\beta,\varepsilon_{2,t}=\gamma^{t},\varepsilon_{3,t}=\delta t,$ for $\beta,\delta>0$ and $0<\gamma<1$. For the loss $L(m,y)=0.5(\Phi(m)-\Phi(y))$, we get by independence that $2\mu_{ij,t}=\mathbb{E}(\Phi(m_{i,t}))-\mathbb{E}(\Phi(m_{j,t}))=\Phi\left(\varepsilon_{i,t}/\sqrt{2}\right)-\Phi\left(\varepsilon_{j,t}/\sqrt{2}\right)$ for $t\in\mathbb{N}$, $i,j=1,2,3$, thus
@@
We let $\beta=0.6,\gamma=0.998$ and $\delta=0.008$. Then, one can numerically show that $\mathcal{M}^{\textrm{w},\star}_{t}=\{3\}$ for $t\leq 153$, $\mathcal{M}^{\textrm{w},\star}_{t}=\{1\}$ for $153<t<550$, and $\mathcal{M}^{\textrm{w},\star}_{t}=\{2\}$ for $t\geq 550$.
@@
Again, we show in Supplement D how one can reduce the average size of the SMCS under a weaker coverage guarantee.
@@ 5.1.3 Results
Throughout the paper we have used a significance level equal to $\alpha=0.1$, but clearly there is freedom to do otherwise. The choice of $\alpha$ affects the interpretation of the SMCS. For example, when $\alpha=0.1$, we can deduce that the SMCS contains the optimal Covid-19 forecasting model with 90% confidence.
The choice of $\alpha$ also affects the size of the SMCS. Larger values of $\alpha$ would lead to earlier detections and to a smaller SMCS. However, they would also translate to a lower level of confidence. In practice, we should keep this tradeoff in mind when choosing $\alpha$. Our goal to narrow down the original set of models to a small SMCS competes with the high level of confidence that is necessary in various applications.
@@ 6 Discussion
We have provided coverage guarantees of all superior models with high probability uniformly over time. Other coverage guarantees could be of interest but require somewhat different approaches. The literature on inference on the argmin of a vector of means typically focuses on the guarantee that each superior model is covered with high probability. Future work could consider time-uniform extensions of such guarantees. Furthermore, one could aim for a time-uniform bound of the expected fraction of wrongly rejected models by the total number of rejected models. This can be achieved by time-uniform control of the false discovery rate and we give first results in this direction in Supplement E.
We have mainly focused on (conditionally) bounded score differences. However, important loss functions such as the squared error or the logarithmic score do not satisfy this requirement. Nevertheless, our general theory is not limited to bounded score differences. We have provided a simulation study for SMCSs with the squared error loss in Supplement G.1. Nonetheless, further research is warranted to extend the understanding and implementation of SMCSs to encompass more general loss functions. Finally, it would be interesting to examine how our results extend to important information criteria, and to apply sequential model confidence sets for sequential model selection.
@@ Acknowledgments
The comments of three anonymous referees significantly improved the quality of the paper. We thank Lutz Dümbgen, Peter Grünwald, and Tim Stephan for valuable discussions and inputs. Sebastian Arnold and Johanna Ziegel gratefully acknowledge financial support from the Swiss National Science Foundation. Computations for the simulations and parts of the wind gust case study have been performed on UBELIX (https://ubelix.unibe.ch/), the HPC cluster of the University of Bern. Benedikt Schulz gratefully acknowledges funding within the project C5 “Dynamical feature-based ensemble postprocessing of wind gusts within European winter storms” of the Transregional Collaborative Research Center SFB/TRR 165 “Waves to Weather” funded by the German Research Foundation (DFG). We thank Cristina Primo, Reinhold Hess and Sebastian Trepte for providing the forecast and observation data, and Robert Redl for assistance in data handling.
@@ Appendix B Conditionally bounded score differences
Lemma B.2.
For $F_{1},F_{2}\in\mathcal{P}(\mathbb{R})$, the CRPS difference $\delta(y)=\textrm{CRPS}(F_{1},y)-\textrm{CRPS}(F_{2},y)$ is bounded in $y\in\mathbb{R}$, and attains its minimum and maximum at a crossing point of $F_{1}$ and $F_{2}$ (including $\{-\infty,+\infty\})$, respectively.
Proof.
We have $\delta(y)=\int F_{1}(x)^{2}-F_{2}(x)^{2}\,\mathrm{d}x+2\int_{y}^{\infty}F_{2}(x)-F_{1}(x)\,\mathrm{d}x,$ where the first summand is independent of $y\in\mathbb{R}$. The function $x\mapsto F_{2}(x)-F_{1}(x)$ is bounded in [1,1][-1,1], with limit 0 as $x\to\pm\infty$ and changes sign at the crossing points of $F_{1}$ and $F_{2}$. ∎
Remark 4.
The minimum and maximum in Lemma B.2 is practically computable in application as illustrated below with the wind gust case study from Section 5.2. Here, the forecasts are either given as parametric (truncated logistic) distributions (EMOS, EMOS-GB, DRN), as distributions with finite support (ENS, MBM, IDR, HEN), or reported as quantiles on a fine grid of quantile levels (QRF, BQN), where the latter two groups may be considered as the same from a computational point of view. Thus, two CDFs may have infinitely many crossing points if they are either identical or if two (discrete) distribution are equal on some given interval. In the former case, the loss difference is trivially equal zero, and in the later case it is straightforward to check that, on the given interval, the absolute value of the CRPS difference is maximized at the boundaries of the interval. In either case, it is thus sufficient to consider a finite number of crossing points. Moreover, an analysis across all stations shows that the values of the maximum bounds are generally around 10, and only rarely reach about 20. Typically, we get these maximum bounds if we compare ENS with some other model, since the ENS forecasts’ relatively large jumps result in big CRPS differences (area between the CDFs) when comparing them to (approximately) smooth distributional forecasts.
@@
In the spirit of Vovk and Wang (2021), we say that non-negative stochastic processes $(E_{1,t}^{\star})_{t\in\mathbb{N}},\dots,(E_{m,t}^{\star})_{t\in\mathbb{N}}$ for the hypotheses $\mathcal{H}_{1},\dots,\mathcal{H}_{m}$ are family-wise valid (FWV) if there exists a dominating family of e-processes, that is, if there exists a family $\{(E^{\mathbb{Q}}_{t})_{t\in\mathbb{N}}\mid\mathbb{Q}\in\mathfrak{B}(\Omega)\}$ such that $(E^{\mathbb{Q}}_{t})_{t\in\mathbb{N}}$ is an e-process for $\{\mathbb{Q}\}$ for all $\mathbb{Q}\in\mathfrak{B}(\Omega)$, and for any $i=1,\dots,m$ with $\mathbb{Q}\in\mathcal{H}_{i}$, it holds that $E_{t}^{\mathbb{Q}}\geq E_{i,t}^{\star}$. Note that each process $(E^{\star}_{i,t})_{t\in\mathbb{N}}$ is an e-process for $\mathcal{H}_{i}$ by monotonicity.
@@
Suppose that we adjust the e-processes $(E_{1,t})_{t\in\mathbb{N}},\dots,(E_{m,t})_{t\in\mathbb{N}}$ by the closure principle with respect to the same e-merging function $f$ for all $t\in\mathbb{N}$. Then it follows that $(E^{\star}_{1,t})_{t\in\mathbb{N}},\dots,(E^{\star}_{m,t})_{t\in\mathbb{N}}$ are family-wise valid e-processes for the hypotheses $\mathcal{H}_{1},\dots,\mathcal{H}_{m}$, since $(E_{t}^{\mathbb{Q}})_{t\in\mathbb{N}}$ are e-processes for $\{\mathbb{Q}\}$ for all $\mathbb{Q}\in\mathfrak{B}(\Omega)$. To see this, let $\tau$ be a stopping time. Then, $\mathbb{E}_{\mathbb{Q}}(E_{\tau}^{\mathbb{Q}})=\mathbb{E}_{\mathbb{Q}}(f(E_{i,\tau},i\in I_{\mathbb{Q}}))\leq 1$, since $E_{i,\tau}$, $i\in I_{\mathbb{Q}}$ are e-values under $\mathbb{Q}$.
@@
where $E_{t}^{\mathbb{Q}}=f(E_{i\cdot,t},i\in I_{\mathbb{Q}})$ with $I_{\mathbb{Q}}=\{i\mid\mathbb{Q}\in\mathcal{H}^{\bullet}_{i\cdot}\}$. It follows directly that the running intersection $\widetilde{\mathcal{M}}_{t}=\bigcap_{r\leq t}\widehat{\mathcal{M}}_{r},t\in\mathbb{N}$, is an SMCS as well. ∎
@@
$\widehat{\mathcal{M}}_{t}=\left\{i\in\mathcal{M}_{0}\mid E_{ij,t}^{\star}<1/\alpha,\text{ for all }j\in\mathcal{M}_{0}\right\},\quad t\in\mathbb{N}.$
@@
Appendix D SMCSs for marginal coverage
In this section, we illustrate how our methods readily extend to the weaker marginal coverage guarantee where the multiple testing correction from the previous section may be omitted.
In analogy to Section 2, we call we call a sequence $(\check{\mathcal{M}}_{t})_{t\in\mathbb{N}}\subseteq\mathcal{M}_{0}$ a marginal coverage SMCS for the superior objects $\mathcal{M}^{\bullet,\star}$, $\bullet\in\{\mathrm{s},\mathrm{uw}\}$ at level $\alpha\in(0,1)$ if, for any $\mathbb{Q}\in\mathfrak{B}(\Omega)$,
$\mathbb{Q}(\forall t\geq 1:i\in\check{\mathcal{M}}_{t})\geq 1-\alpha\quad\textrm{for all }i\in\mathcal{M}^{\bullet,\star}.$ (15)
We refer to condition (15) as time-uniform marginal coverage. Clearly, this is a weaker requirement than given at (4), forming a time-uniform simultaneous coverage.
As in Section 3.2, for $\bullet\in\{\mathrm{s},\mathrm{uw}\}$, assume that, for any $i,j\in\mathcal{M}_{0}$, $(E_{ij,t})_{t\in\mathbb{N}}$ is an e-process for the hypothesis $\mathcal{H}^{\bullet}_{ij}$, and let $E_{i\cdot,t}=1/(m-1)\sum_{j\neq i}E_{ij,t}$ be the arithmetic mean e-process for the intersection hypothesis $\mathcal{H}^{\bullet}_{i\cdot}=\cap_{j\neq i}\mathcal{H}^{\bullet}_{ij}$. For $\alpha\in(0,1)$, define
$\check{\mathcal{M}}_{t}=\left\{i\in\mathcal{M}_{0}\mid E_{i\cdot,t}<1/\alpha\right\},\quad t\in\mathbb{N}.$ (16)
Then, the following adaption of Theorem 3.1 follows directly.
Theorem D.1.
For any $\alpha\in(0,1)$, the sequence $(\check{\mathcal{M}}_{t})_{t\in\mathbb{N}}$ defined at (16) is a marginal coverage SMCS at level $\alpha$ for $\mathcal{M}^{\bullet,\star}$, $\bullet\in\{\mathrm{s},\mathrm{uw}\}$, and so is its running intersection.
Remark 6.
If $|\mathcal{M}^{\bullet,\star}|=1$, $\bullet\in\{\mathrm{s},\mathrm{uw}\}$, then (4) and (15) coincide, and we do not have to correct for multiple testing if there exists exactly one superior object.
Since weakly superior objects can change over time we have to adapt the definition of marginal coverage at (15) for weakly superior objects. For each $i\in\mathcal{M}_{0}$, we let
$I^{\star}(i)=\{t\in\mathbb{N}\mid\Delta_{ij,t}\leq 0\textrm{ a.s. for all }j\in\mathcal{M}_{0}\},$
and call a sequence $(\tilde{\mathcal{M}}_{t})_{t\in\mathbb{N}}\subseteq\mathcal{M}_{0}$ a marginal coverage SMCS for the weakly superior objects $(\mathcal{M}^{\textrm{w},\star}_{t})_{t\in\mathbb{N}}$, at level $\alpha\in(0,1)$ if, for any $\mathbb{Q}\in\mathfrak{B}(\Omega)$,
$\mathbb{Q}(\forall t\in I^{\star}(i):i\in\tilde{\mathcal{M}}_{t})\geq 1-\alpha\quad\textrm{for all }i\in\mathcal{M}_{0}.$
Next, we show how one can construct marginal coverage SMCSs for the weakly superior objects. For $i\in\mathcal{M}_{0}$, define $\boldsymbol{\Delta}_{i,t}=(\Delta_{ij,t})_{j=1}^{m}$ and, for $\boldsymbol{X}=(x_{i})_{i=1}^{m}\in\mathbb{R}^{m}$,
$M_{t}^{(i)}(\boldsymbol{X})=\frac{1}{m}\sum_{j=1}^{m}M_{ij,t}(x_{j}),$
where $M_{ij,t}(x)$ is such that $(M_{ij,t}(\Delta_{ij,t}))_{t\in\mathbb{N}}$ is a test supermartingale for any $\mathbb{Q}\in\mathfrak{B}(\Omega)$. Then, $(M_{t}^{(i)}(\boldsymbol{\Delta}_{i,t}))_{t\in\mathbb{N}}$ is a test supermartingale for any $\mathbb{Q}\in\mathfrak{B}(\Omega)$. For $\alpha\in(0,1)$, we define $m$ confidence regions
$C_{t}^{(i)}=\left\{\boldsymbol{X}\in\mathbb{R}^{m}\mid M_{t}^{(i)}(\boldsymbol{X})<\frac{1}{\alpha}\right\},\quad i\in\mathcal{M}_{0},$
and set
$\tilde{\mathcal{M}}_{t}=\{i\in\mathcal{M}_{0}\mid\forall j\not=i:C_{t}^{(i)}\cap\mathbb{R}_{j,-}^{m}\not$$\emptyset\}$, (17)
where $\mathbb{R}_{j,-}^{m}=\{\boldsymbol{X}\in\mathbb{R}^{m}\mid x_{j}\leq 0\}$. This means that, analogously to Section 3.3, we exclude $i$ from $\tilde{\mathcal{M}}_{t}$ if there exists a $j\not=i$ such that we may reject $\Delta_{ij,t}\leq 0$, which is the case, if and only if $C_{t}^{(i)}\cap\mathbb{R}_{j,-}^{m}=\emptyset$ for some $j\not=i$.
Theorem D.2.
For any $\alpha\in(0,1)$, the sequence $(\tilde{\mathcal{M}}_{t})_{t\in\mathbb{N}}$ defined at (17) is a marginal coverage SMCS for the weakly superior objects $(\mathcal{M}^{\textrm{w},\star}_{t})_{t\in\mathbb{N}}$.
Proof.
For any $\mathbb{Q}\in\mathfrak{B}(\Omega)$, and any $i\in\mathcal{M}_{0}$, we have
$\mathbb{Q}(\exists t\in I^{\star}(i):i\notin\tilde{\mathcal{M}}_{t})$ $=\mathbb{Q}(\exists t\in I^{\star}(i),\exists j\in\mathcal{M}_{0}:C_{t}^{(i)}\cap\mathbb{R}_{j,-}^{m}=\emptyset)$
$\leq\mathbb{Q}(\exists t\in I^{\star}(i):M_{t}^{(i)}(\boldsymbol{\Delta}_{i,t})\geq 1/\alpha)\leq\alpha\,.$
In Simulations 1 and 2, there is always an unique superior object, so marginal and uniform coverage are equivalent. Therefore, we can compute the SMCS without having to adjust the e-processes. Figure 7 shows that this results in smaller confidence sets. Analogously, we compare the marginal coverage SMCS for the weakly superior objects given by (17) with the original SMCS from Section 3.3. This comparison is illustrated in Figure 8. As expected, the average size of the SMCS given by (17) is generally smaller.
Figure 7: The SMCS shrinks slightly when uniform coverage is replaced by marginal coverage. In Simulations 1 and 2, marginal coverage and uniform coverage are equivalent, due to the existence of a unique superior object.
Figure 8: Size of the SMCS over time in Simulation 3, averaged over $100$ simulations, for marginal coverage (blue) compared to the original approach from Section 3.3 (red), see the original Figure 2. Also for the marginal coverage SMCS we have 100% coverage of the superior object at all time points.
Appendix E SMCSs by FDR control